OpenStudy (anonymous):

Solve C = 2πr (solve for "r") C = 15

OpenStudy (anonymous):

C = 2πr to solve for r, divide both sides by 2π C/2π = r

OpenStudy (anonymous):

This is the second part to the question: b. If a circle's circumference is 15 inches, what is its radius? Leave the symbol π in your answer. The radius of the circle is _______

OpenStudy (petiteme):

divide both sides by 2π C/2π = r just substitute 15 to C :) then divide 2π :)

OpenStudy (anonymous):

15=2(pi)r 7.5=(pi)r 7.5/(pi)=r

OpenStudy (anonymous):

would i use 3.14 or 22/7?

OpenStudy (petiteme):

3.14 :)

OpenStudy (anonymous):

anything else

OpenStudy (anonymous):

so if I did my work right, "r" would equal 2.39?

OpenStudy (petiteme):

YUP! :)

OpenStudy (anonymous):

awesome! thanks you guys!

OpenStudy (petiteme):

You're welcome! Anytime :)